# IMU: Influence-guided Machine Unlearning

### v2

@@
Abstract
Recent studies have shown that deep learning models are vulnerable to attacks and tend to memorize training data points, raising significant concerns about privacy leakage. This motivates the development of machine unlearning (MU), i.e., a paradigm that enables models to selectively forget specific data points upon request. However, most existing MU algorithms require partial or full fine-tuning on the retain set. This necessitates continued access to the original training data, which is often impractical due to privacy concerns and storage constraints. A few retain-data-free MU methods have been proposed, but some rely on access to auxiliary data and precomputed statistics of the retain set, while others scale poorly when forgetting larger portions of data. In this paper, we propose Influence-guided Machine Unlearning (IMU), a simple yet effective method that conducts MU using only the forget set. Specifically, IMU employs gradient ascent and innovatively introduces dynamic allocation of unlearning intensities across different data points based on their influences. This adaptive strategy significantly enhances unlearning effectiveness while maintaining model utility. Results across vision and language tasks demonstrate that IMU consistently outperforms existing retain-data-free MU methods.
1 Introduction
@@
Deep learning (DL) models are widely deployed in various applications, but their vulnerability to adversarial attacks, such as membership inference attacks (Rezaei and Liu 2021; Watson et al. 2022) and model inversion attacks (Zhu, Liu, and Han 2019; Geiping et al. 2020; Balunović et al. 2022), raises significant concerns about privacy leakage. In response, legislation such as GDPR grants users the right to be forgotten, compelling models to remove data points upon request. The straightforward method is to retrain the model without using the forgetting data from scratch. However, this is highly inefficient, particularly for large-scale models trained on massive datasets. As a result, machine unlearning (MU), i.e., removing the influence of specific data points from a well-trained model while preserving the model utility on unrelated information without the need for costly and prolonged retraining, has emerged as a crucial approach for safeguarding data privacy.
Current MU methods typically require fine-tuning on the retain data, which are part of the original training data. However, accessing the original training data is often impractical in real-world deployment. Recently, a few studies have sought to overcome this limitation by developing retain-data-free MU methods (Cha et al. 2024; Bonato, Cotogni, and Sabetta 2024; Foster et al. 2025). Among these efforts, Cha et al. (2024) presents instance-wise unlearning that generates adversarial examples w.r.t. the forgetting data and fine-tunes over these examples, specifically targeting model parameters responsible for the correct classification of the forgetting data points. Meanwhile, Bonato, Cotogni, and Sabetta (2024) introduces selective-distillation for class and architecture-agnostic Learning (SCAR), which uses the Mahalanobis distance to shift the feature vectors w.r.t. the forgetting data toward the nearest wrong class distribution, then distills the knowledge from the original model into the scrubbed model using out-of-distribution images.
While these retain-data-free methods have shown promising results in deleting data from DL models, they exhibit important limitations. Cha et al. (2024) only consider evaluation when the forgetting data is randomly selected from the training data (e.g., around 0.5% on CIFAR-10), raising concerns about its scalability to larger forgetting sets. SCAR needs an auxiliary dataset to maintain model performance and assumes that the statistics (e.g., mean, and covariance) of the retain set are stored, which may not always be available in practical scenarios. These constraints highlight the need for a more robust and generalizable retain-data-free unlearning framework that can handle diverse forgetting scenarios without external data requirements.
Influence function (Koh and Liang 2017), which estimates parameter changes induced by data point removal without full retraining, has attracted substantial attention and research interest in recent years (Basu, Pope, and Feizi 2021; Bae et al. 2022; Chhabra et al. 2024). While influence-based methods have been adapted for MU (Sekhari et al. 2021; Neel, Roth, and Sharifi-Malvajerdi 2021; Mehta et al. 2022; Wu, Hashemi, and Srinivasa 2022), two critical challenges persist: (1) Influence function approximations are known to be fragile in deep learning (Basu, Pope, and Feizi 2021). The influence estimation is fairly accurate for shallow networks where the loss function is convex, but prone to error in deep models due to the non-convexity of their loss landscapes; (2) for large neural networks with complex architectures and millions of parameters, computing inverse Hessian-vector products, even with the Fisher Information approximation to estimate the influence function, remains prohibitively expensive (Mehta et al. 2022).
In this work, we argue that the full potential of the influence function on MU remains underexplored. While conventional influence-based unlearning (IU) methods primarily focus on computational efficiency (Mehta et al. 2022), they overlook two critical aspects: (1) the stability of influence estimates in deep non-convex networks, and (2) the heterogeneous influence distribution across forgetting data points, where certain samples exert disproportionately stronger effects on the model than others. Therefore, we propose a novel Influence-guided Machine Unlearning (IMU) method to dynamically allocate different attention to the forgetting data points based on their influence value. We denote the final fully connected layer as the classifier and the preceding layers as the feature extractor. To improve the accuracy of the estimation of influence value and avoid expensive matrix inversion computations, we view the feature representations extracted by the feature extractor as the input, and then estimate the influence value at the classifier. In this way, the classifier is convex, and the influence estimates would be more reliable. Experimental results demonstrate that our proposed algorithm effectively removes the influence of forgetting data while preserving model utility, showing both efficiency and stability in the MU process.
Our main contributions are summarized as:
  • We propose a retain-data-free MU method, IMU, which leverages influence functions on non-convex models by estimating the influence at the final fully connected layer. This design avoids full-model Hessian inversion while maintaining estimation accuracy.
  • IMU automatically adjusts the unlearning strength for each forgetting data point proportionally to its influence score, applying more aggressive parameter updates for high-influence samples while preserving knowledge from less influential ones.
@@
  • Extensive experiments across vision and language tasks show that our proposed method IMU consistently surpasses state-of-the-art retain-data-free MU methods in both forget quality and model utility.
2 Related Work
2.1 Machine Unlearning
Figure 1: Examples of data points with different degrees of influence on the data set that belong to their own class.
@@ 2.1 Machine Unlearning
Current MU algorithms typically fall into two categories: exact unlearning methods (Ginart et al. 2019; Cao and Yang 2015; Romero, Barrio, and Belanche 2007; Karasuyama and Takeuchi 2010) that use data partitioning strategies (Bourtoule et al. 2021) or only focus on traditional machine learning models like $k$-means clustering (Ginart et al. 2019), and approximate unlearning methods (Golatkar, Achille, and Soatto 2020a; Chen et al. 2023; Heng and Soh 2023; Kumari et al. 2023; Kurmanji et al. 2023; Foster, Schoepf, and Brintrup 2024; Lyu et al. 2024; Bui et al. 2024; Ko et al. 2024; Lin et al. 2024; Spartalis et al. 2025; Alberti et al. 2025; Wu et al. 2025; Wu and Harandi 2024a) that trade perfect deletion for computational efficiency. So far, these MU methods have demonstrated effectiveness through two key metrics: (1) degraded performance on the forgetting data, and (2) maintained utility on retain and unseen set.
@@
Gradient ascent (GA)-based MU methods (Wu, Dobriban, and Davidson 2020; Gandikota et al. 2023; Jang et al. 2022) perform unlearning by moving model parameters away from the forgetting data through gradient ascent. While simple to implement, these methods can be sensitive to hyperparameters and may interfere with retain knowledge (Jia et al. 2023; Fan et al. 2024a). Fisher unlearning methods (Golatkar, Achille, and Soatto 2020a, b; Golatkar et al. 2021) assume that the unlearned model and the retrained model are close to each other in parameter space, and formulate Fisher Forgetting, which induces unlearning by injecting noise into the parameters proportional to their relative importance in the forgetting data compared to the retain set. Influence function-based methods estimate the impact of individual data points to perform approximate unlearning. Guo et al. (2020) propose a one‐step Newton‐update procedure based on the influence function to excise the effect of specific data points from a pretrained model. Subsequent works (Mehta et al. 2022; Liu et al. 2022) focus on ways for more efficient computation over the estimation. However, these methods make strong convexity assumptions and are still computationally expensive for large-scale models (Mehta et al. 2022). Two-stage methods (Jia et al. 2023; Fan et al. 2024b; Wu and Harandi 2024b) have emerged as a practical alternative, first erasing information related to the forgetting data, then fine-tuning on retain data. These methods avoid expensive computations while achieving strong empirical results, particularly when combined with sparsity techniques.
These MU methods, however, still require access to the retain set, presenting practical challenges when storage constraints or privacy concerns limit data availability. Recent work has begun addressing this limitation through alternative approaches (Cha et al. 2024; Bonato, Cotogni, and Sabetta 2024; Foster et al. 2025). Cha et al. (2024) proposes an instance-wise unlearning method that generates adversarial examples w.r.t. the forgetting data and fine-tunes the model on these examples, specifically targeting parameters responsible for the correct classification of the forgetting data. Meanwhile, Bonato, Cotogni, and Sabetta (2024) introduces SCAR, which leverages the Mahalanobis distance to shift feature representations of forgetting data toward the nearest incorrect class distribution. The scrubbed model then undergoes knowledge distillation using out-of-distribution images to preserve model utility. NPO (Zhang et al. 2024) treats the forgetting data as negative examples in DPO (Rafailov et al. 2023), presenting an adaptive control in large language model (LLM) unlearning. Fan et al. (2024a) further improve NPO via a reference-free method SimNPO (Meng, Xia, and Chen 2024), which addresses the reference model bias issue.
These advances are particularly valuable for real-world deployment, where retaining full training data is often impractical. However, existing retain-data-free MU methods exhibit notable limitations. Cha et al. (2024) evaluates only on small, randomly selected forgetting sets, raising concerns about scalability to larger deletions. Bonato, Cotogni, and Sabetta (2024) relies on an auxiliary dataset to preserve model utility, and requires that the statistics (e.g., mean and covariance) of retain data are stored, a requirement that may not always be feasible. These constraints underscore the need for a more robust and generalizable retain-free unlearning framework capable of handling diverse forgetting scenarios without external data dependencies.
2.2 Influence Function
Influence functions (Hampel 1974; Koh and Liang 2017; Pruthi et al. 2020) assess how infinitesimal changes to training data weights affect performance on a validation set, by analyzing their impact on a target evaluation metric.
@@ 2.2 Influence Function
For linear models, influence functions are well-defined thanks to the convexity of the loss function, but in deep learning, where loss functions are typically non-convex, the behavior of influence functions remains poorly understood (Basu, Pope, and Feizi 2021). Basu, Pope, and Feizi (2021) presents a comprehensive empirical study on the reliability of influence functions in DNNs. The findings reveal that factors such as network architecture, depth, width, parameterization, and regularization significantly impact influence estimation accuracy. Overall, the results demonstrate that influence functions in DNNs frequently fail to accurately predict the effects of retraining, leading to the conclusion that influence estimates are often brittle and unreliable.
However, recent works Bae et al. (2022); Epifano et al. (2023) have challenged these claims, making their applicability in DNNs an ongoing topic of debate. So far, several strategies have been proposed to extend influence functions to non-convex models: (1) replacing non-convex embeddings with those from linear models (Li and Liu 2022; Chhabra et al. 2024); (2) adding damping terms to make the Hessian matrix positive definite (Koh and Liang 2017; Han, Wallace, and Tsvetkov 2020); and (3) deriving task-specific or second-order influence functions (Basu, You, and Feizi 2020; Alaa and Van Der Schaar 2020).
Existing influence function-based MU methods mainly focus on computational efficiency (Mehta et al. 2022; Jia et al. 2023), they overlook two key challenges: (1) the instability of influence estimates in deep, non-convex models, and (2) the heterogeneous influence distribution across the forgetting data points, where certain samples have a disproportionately large impact on the model. In response, we aim to enhance influence-based unlearning by addressing these challenges through a more stable and adaptive influence estimation framework.
3 Methodology
@@
In this section, we propose IMU, a retain-data-free unlearning framework that scrubs data from a model by dynamically eliminating its influence from the model. Throughout the paper, we denote scalars and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, $\bm{a}$}, and $\bm{A}$}).
3.1 Preliminaries
@@
Influence Function. Refer to (Koh and Liang 2017), denote the empirical risk minimizer of the loss over the training set as ${\bm{\theta}}^{\ast}\coloneqq\operatorname*{arg\,min}_{{\bm{\theta}}}\frac{1}{N}\sum_{i=1}^{N}\ell({\bm{x}}_{i},{\bm{y}}_{i};{\bm{\theta}})$, the change in model parameters caused by an infinitesimal up-weighting of the training data point $\bm{x}$} can be approximated by
${\mathcal{I}}({\bm{x}})\coloneqq-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big{)},$ (1)
where ${\bm{H}}_{{\bm{\theta}}^{\ast}}=\frac{1}{N}\sum_{i=1}^{N}\nabla^{2}_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}}_{i})\big{)}$. Then, the change in the loss value for ${\mathcal{D}}_{t}$ when the training data point $\bm{x}$} is removed from the training set can be approximated by
${\mathcal{I}}({\bm{x}},{\mathcal{D}}_{t})\coloneqq-\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\mathcal{D}}_{t})\big{)}^{\top}{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big{(}f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big{)}.$ (2)
@@
Proofs can be found in  (Koh and Liang 2017). However, the derivation relies on the assumption that the underlying loss function is strictly convex w.r.t. the model parameters and that the Hessian matrix is positive definite. Besides, for deep neural networks, computing the exact inverse of the Hessian matrix is computationally expensive (Mehta et al. 2022).
3.2 IMU
Our proposed method, IMU, addresses two key challenges in applying influence functions to deep neural networks: (1) the instability of influence estimates in highly non-convex networks, and (2) the heterogeneous influence distribution among forgetting samples, where certain data points have disproportionately large effects on the model.

### v3

@@
Abstract.
Machine Unlearning (MU) aims to selectively erase the influence of specific data points from pretrained models. However, most existing MU methods rely on the retain set to preserve model utility, which is often impractical due to privacy restrictions and storage constraints. While several retain-data-free methods attempt to bypass this using geometric feature shifts or auxiliary statistics, they typically treat forgetting samples uniformly, overlooking their heterogeneous contributions. To address this, we propose Influence-guided Machine Unlearning (IMU), a principled method that conducts MU using only the forget set. Departing from uniform Gradient Ascent (GA) or implicit weighting mechanisms, IMU leverages influence functions as an explicit priority signal to allocate unlearning strength. To circumvent the prohibitive cost of full-model Hessian inversion, we introduce a theoretically grounded classifier-level influence approximation. This efficient design allows IMU to dynamically reweight unlearning updates, aggressively targeting samples that most strongly support the forgetting objective while minimizing unnecessary perturbation to retained knowledge. Extensive experiments across vision and language tasks show that IMU achieves highly competitive results. Compared to standard uniform GA, IMU maintains identical unlearning depth while enhancing model utility by an average of 30%, effectively overcoming the inherent utility-forgetting trade-off.
Machine unlearning, Retain-data-free, Influence function
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: ; ; isbn: 978-1-4503-XXXX-X/2026/06
1. Introduction
Figure 1. Visualizing the heterogeneous influence of the forgetting data. (Left) The scatter plot reveals a wide variance in per-sample influence values, confirming that individual data points contribute non-uniformly to the learned model. (Right) Example images categorized by their influence magnitudes.
@@ 1. Introduction
Deep learning (DL) models are widely deployed in various applications, but their vulnerability to adversarial attacks, such as membership inference attacks (Rezaei and Liu, 2021) and model inversion attacks (Zhu et al., 2019; Geiping et al., 2020; Balunović et al., 2022), raises significant concerns about privacy leakage. In response, legislation such as GDPR (Voigt and Von dem Bussche, 2017) grants users the right to be forgotten, compelling models to remove data points upon request. Retraining from scratch is often computationally prohibitive for modern large-scale models, motivating the development of machine unlearning (MU), i.e., a paradigm aimed at efficiently removing the effect of the forgetting data while preserving utility on retain data.
Current MU methods typically require fine-tuning on the retain data, which are part of the original training data. However, this retain data may be entirely unavailable due to privacy restrictions, ownership, or storage constraints in many real-world deployments. Recently, a few studies have sought to overcome this limitation by developing retain-data-free MU methods (Cha et al., 2024; Bonato et al., 2024; Foster et al., 2025). While these methods show promising unlearning capabilities, they often operate under restrictive assumptions (e.g., requiring access to auxiliary data or stored dataset statistics).
More importantly, a key challenge that remains underexplored is the heterogeneous contribution of individual forgetting data points. In practice, forgetting data points exhibit highly non-uniform influence on the learned model: a small subset may contribute dominantly to the model’s fit on the forget set ${\mathcal{D}}_{f}$, while others provide marginal or redundant contributions, and may even introduce unnecessary drift to utility-relevant parameters when treated uniformly during unlearning. Figure 1 illustrates this heterogeneity when conducting class-wise and sample-wise unlearning in CIFAR-10, where per-sample influence values span a wide range. Samples with high removal influence (i.e., whose removal would significantly raise the loss on ${\mathcal{D}}_{f}$) represent data points that the model has fitted most deeply; they therefore require the strongest unlearning effort to effectively erase. In contrast, low-influence samples contribute little to the model’s knowledge of ${\mathcal{D}}_{f}$, and aggressively unlearning them provides diminishing returns while risking unnecessary perturbation to retained knowledge.
To address this heterogeneous nature, we propose Influence-guided Machine Unlearning (IMU), a retain-data-free framework that dynamically allocates unlearning strength based on each sample’s true contribution. Instead of relying on standard uniform Gradient Ascent (GA), which blindly updates parameters and risks catastrophic utility collapse, we leverage influence functions (Koh and Liang, 2017) as an explicit priority signal. By quantifying how removing a specific data point affects the forgetting objective, IMU directs the limited unlearning budget to where it is most needed. It aggressively erases high-influence samples while safeguarding the model from unnecessary drift caused by less informative ones.
However, directly applying influence functions to deep neural networks presents significant practical challenges. Theoretical influence estimation relies on convex loss landscapes, and computing the required inverse Hessian-vector products is computationally prohibitive for modern models with millions of parameters (Basu et al., 2021; Mehta et al., 2022). To overcome these computational bottlenecks, we introduce a theoretically grounded classifier-level influence approximation. By viewing the network as a fixed feature extractor followed by a shallow prediction head, we restrict the Hessian computation exclusively to the classifier. This formulation not only yields a convex subproblem for reliable estimation but also drastically reduces the computational overhead. Ultimately, IMU provides a principled, highly efficient approach to retain-data-free MU. Experimental results demonstrate that our proposed algorithm effectively removes the influence of forgetting data while preserving model utility, consistently outperforming existing retain-data-free baselines.
Our contributions are as follows:
  • We propose IMU, a principled retain-data-free MU method that explicitly addresses the heterogeneous contributions of forgetting data. By utilizing influence scores as a priority signal, IMU dynamically allocates unlearning strength, i.e., aggressively targeting high-influence samples while safeguarding retained knowledge from less informative ones.
  • To enable the efficient use of influence functions within this framework, we introduce a theoretically grounded classifier-level influence approximation. This design successfully circumvents the intractable cost of full-model Hessian inversion on deep non-convex models, maintaining high estimation fidelity with minimal overhead.
@@
  • Extensive experiments across vision and language tasks show that IMU consistently outperforms existing retain-data-free MU methods, achieving a comparable balance between unlearning efficacy and model utility preservation.
2. Related Work
2.1. Machine Unlearning
@@
MU algorithms typically fall into two categories: exact MU methods (Ginart et al., 2019; Cao and Yang, 2015; Bourtoule et al., 2021) that use data partitioning strategies or only focus on traditional machine learning models like $k$-means clustering (Ginart et al., 2019), and approximate MU methods that trade perfect deletion for computational efficiency. The latter include parameter perturbation (Golatkar et al., 2020a; Chen et al., 2023; Kumari et al., 2023; Foster et al., 2024a), incremental strategies (Heng and Soh, 2023; Kurmanji et al., 2023; Lyu et al., 2024; Bui et al., 2024), and sparsity or data-attribution techniques (Ko et al., 2024; Lin et al., 2024; Spartalis et al., 2025; Alberti et al., 2025; Wu et al., 2025; Foster et al., 2024b; Goel et al., 2023). So far, these MU methods have demonstrated effectiveness through two key metrics: (1) degraded performance on the forgetting data, and (2) maintained utility on retain and unseen set.
@@
Gradient ascent (GA)-based MU (Wu et al., 2020; Gandikota et al., 2023; Jang et al., 2022) perform unlearning by moving model parameters away from the forgetting data through gradient ascent. While simple to implement, in practice, these methods require control mechanisms to prevent damaging the model’s utility (Jia et al., 2023; Fan et al., 2024a). Fisher unlearning (Golatkar et al., 2020a, b, 2021) deem the unlearned model and the retrained model are close to each other in parameter space, and formulate Fisher Forgetting, which induces unlearning by injecting noise into the parameters proportional to their relative importance in the forgetting data compared to the retain set. Influence function-based methods estimate the impact of individual data points to perform approximate unlearning. Guo et al. (2020) propose a one‐step Newton‐update procedure based on the influence function to excise the effect of specific data points from a pretrained model. Subsequent works (Mehta et al., 2022; Liu et al., 2022; Jia et al., 2023) focus on ways for more efficient computation over the estimation. Two-stage methods (Jia et al., 2023; Fan et al., 2024b) have emerged as a practical alternative, first erasing information related to the forgetting data, then fine-tuning on retain data. These methods avoid expensive computations while achieving strong empirical results, particularly when combined with sparsity techniques.
Retain-data-free MU. While the aforementioned methods show strong empirical results, they rely on access to the retain set, presenting challenges under strict privacy or storage constraints. Recent studies (Cha et al., 2024; Bonato et al., 2024; Foster et al., 2025) have pioneered retain-data-free methods to address this. Cha et al. (2024) propose instance-wise unlearning by generating and fine-tuning on adversarial examples w.r.t. the forgetting data. Bonato et al. (2024) leverages Mahalanobis distance to shift forgetting representations toward incorrect classes, followed by knowledge distillation with auxiliary images for preserving model utility. In the large language model (LLM) domain, Negative Preference Optimization (NPO) (Zhang et al., 2024; Fan et al., 2024a) treats forgetting data as negative examples within a preference optimization framework.
Despite their promising unlearning capabilities, a critical aspect that remains underexplored across these methods is the heterogeneous contributions of individual forgetting data points. Most existing approaches apply a uniform unlearning objective across the entire forget set. While recent frameworks like NPO implicitly introduce instance-level weighting through preference optimization, they scale updates based on output-level prediction confidence (e.g., log-likelihoods) rather than a direct measure of a sample’s parametric contribution. Consequently, these methods risk over-unlearning low-influence samples or under-unlearning critical ones. Complementing these pioneering efforts, our work explicitly tackles this heterogeneity by introducing a theoretically grounded framework that dynamically allocates unlearning budget based on the precise parametric influence of each data point.
2.2. Influence Function in Deep Learning
@@
Influence functions (Hampel, 1974; Koh and Liang, 2017; Pruthi et al., 2020) assess how infinitesimal changes to training data weights affect performance on a validation set, by analyzing their impact on a target evaluation metric. For linear models, influence functions are well-defined thanks to the convexity of the loss landscape, but in deep learning, where loss landscapes are typically non-convex, the behavior of influence functions remains poorly understood (Basu et al., 2021). Basu et al. (2021) presents a comprehensive empirical study on the reliability of influence functions in deep neural networks (DNNs). The findings reveal that factors such as network architecture, depth, width, parameterization, and regularization significantly impact influence estimation accuracy. Overall, the results demonstrate that influence functions in DNNs frequently struggle to accurately predict the effects of retraining, leading to the conclusion that influence estimates are often brittle and unreliable.
Although recent works (Bae et al., 2022; Epifano et al., 2023) have defended their applicability in DNNs, extending them remains non-trivial. Existing strategies include replacing non-convex embeddings with linear ones (Li and Liu, 2022; Chhabra et al., 2024), adding damping terms to ensure positive-definite Hessians (Koh and Liang, 2017; Han et al., 2020), or deriving second-order formulations (Basu et al., 2020; Alaa and Van Der Schaar, 2020). Furthermore, computing the required inverse Hessian-vector products is computationally prohibitive for modern large-scale models (Mehta et al., 2022). To bypass these limitations, we introduce a theoretically grounded classifier-level approximation, enabling the robust and efficient use of influence functions for MU without requiring full-model Hessian inversion.
3. Methodology
@@
In this section, we propose IMU, a retain-data-free unlearning framework that scrubs data from a model by dynamically eliminating its influence from the model. We consider the unlearning setting where, upon a deletion request, the unlearning procedure is only allowed to access ${\mathcal{D}}_{f}$ and the current model, while ${\mathcal{D}}_{r}$ is unavailable due to privacy, storage, or ownership constraints. Throughout the paper, we denote scalars and vectors/matrices by lowercase and bold symbols, respectively (e.g., $a$, $\bm{a}$}, and $\bm{A}$}).
3.1. Preliminaries
@@
Influence function. Refer to (Koh and Liang, 2017), we denote the empirical risk minimizer of the loss over ${\mathcal{D}}_{\text{train}}$ as ${\bm{\theta}}^{\ast}\triangleq\operatorname*{arg\,min}_{{\bm{\theta}}}\frac{1}{N}\sum_{i=1}^{N}\ell(f_{\bm{\theta}}({\bm{x}}_{i}))$. If upweighting the training data point $\bm{x}$} by some small $\epsilon$, the new optimal parameter is $\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}\triangleq\operatorname*{arg\,min}_{{\bm{\theta}}}\frac{1}{N}\sum_{i=1}^{N}\ell(f_{\bm{\theta}}({\bm{x}}_{i}))+\epsilon\ell(f_{\bm{\theta}}({\bm{x}}))$. The change in model parameters can then be approximated by
(1) ${\mathcal{I}}_{\texttt{up},{\bm{\theta}}^{\ast}}({\bm{x}})\triangleq\left.\frac{d\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}}{d\epsilon}\right|_{\epsilon=0}=-{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big),$
where ${\bm{H}}_{{\bm{\theta}}^{\ast}}=\frac{1}{N}\sum_{i=1}^{N}\nabla^{2}_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}}_{i})\big)$. Consequently, the influence on the loss for ${\mathcal{D}}_{t}$ when upweighting $\bm{x}$} would be:
${\mathcal{I}}_{\texttt{up},\ell}({\bm{x}},{\mathcal{D}}_{t})$ $\triangleq\left.\frac{d\ell\left(f_{\hat{{\bm{\theta}}}_{\epsilon,{\bm{x}}}}({\mathcal{D}}_{t})\right)}{d\epsilon}\right|_{\epsilon=0}$
(2) $=$- $\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\mathcal{D}}_{t})\big)^{\top}{\bm{H}}_{{\bm{\theta}}^{\ast}}^{-1}\nabla_{{\bm{\theta}}}\ell\big(f_{{\bm{\theta}}^{\ast}}({\bm{x}})\big).$
@@
Proofs can be found in  (Koh and Liang, 2017). However, this classical derivation assumes sufficient smoothness and local invertibility of the Hessian, which may not strictly hold for highly non-convex deep networks. Besides, for DNNs, computing the exact inverse of the Hessian matrix is computationally expensive (Mehta et al., 2022).
3.2. IMU
As discussed in § 1, a critical aspect that remains underexplored across retain-data-free MU methods is the heterogeneous contributions of individual forgetting data points. Treating all forgetting samples uniformly (e.g., GA) ignores the fact that certain data points contribute disproportionately to the model’s learned representations. To achieve precise and effective unlearning, it is crucial to dynamically allocate the unlearning budget based on each sample’s actual parametric contribution. While influence functions provide a principled mathematical framework to quantify this sample-wise contribution, directly applying them to DNNs is practically prohibited by the computational cost and instability of high-dimensional Hessians. To bridge this gap, IMU proposes a robust, classifier-level influence estimation strategy to efficiently guide the MU process.